# Deliberate Planning of 3D Bin Packing on Packing Configuration Trees

### v3

@@ Appendix A Implementation Details
$J(\pi)=E_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$ (12)
Where $\gamma\in[0,1]$ is the discount factor, and $\tau=(s_{0},a_{0},s_{1},\ldots)$ is a trajectory sampled based on the policy $\pi$. We extract the feature of state $s_{t}=(\mathcal{T}_{t},n_{t})$ using graph attention networks (Velickovic et al., 2018) for encoding the spatial relations of all space configuration nodes. The context feature is fed to two key components of our pipeline: an actor network and a critic network. The actor network, designed based on a pointer mechanism, weighs the leaf nodes of PCT, which is written as $\pi(a_{t}|s_{t})$. The action $a_{t}$ is an index of selected leaf node $l\in\textbf{L}_{t}$, denoted as $a_{t}=index(l)$. The critic network maps the context feature into a state value prediction $V(s_{t})$, which helps the training of the actor network. The whole network is trained via a composite loss $L=\alpha\cdot L_{actor}+\beta\cdot L_{critic}$ ($\alpha=\beta=1$ in our implementation), which consists of actor loss $L_{actor}$ and critic loss $L_{critic}$. These two loss functions are defined as:
$\begin{array}[]{ll}\displaystyle L_{actor}&\displaystyle=(r_{t}+\gamma V(s_{t+1})-V(s_{t}))\log\pi(a_{t}|s_{t})\\ \displaystyle L_{critic}&\displaystyle=(r_{t}+\gamma V(s_{t+1})-V(s_{t}))^{2}\\ \end{array}$ (13)
Where $r_{t}=c_{r}\cdot w_{t}$ is our reward signal, and we set $\gamma$ as 1 since the packing episode is finite. We adopt a step-wise reward $r_{t}=c_{r}\cdot w_{t}$ once $n_{t}$ is inserted into PCT as an internal node successfully. Otherwise, $r_{t}=0$ and the packing episode ends. The choice of item weight $w_{t}$ depends on the packing preferences. In the general sense, we set $w_{t}$ as the volume occupancy $v_{t}=s^{x}_{n}\cdot s^{y}_{n}\cdot s^{z}_{n}$ of $n_{t}$, and the constant $c_{r}$ is $10/(S^{x}\cdot S^{y}\cdot S^{z})$. For online 3D-BPP with additional packing constraints, this weight can be set as $w_{t}=max(0,v_{t}-c\cdot O(s_{t},a_{t}))$. While the term $v_{t}$ ensures that space utilization is still the primary concern, the objective function $O(s_{t},a_{t})$ guides the agent to satisfy additional constraints like isle friendliness and load balancing. We adopt the ACKTR (Wu et al., 2017) method for training our DRL agent, which iteratively updates an actor and a critic using Kronecker-factored approximate curvature (K-FAC) (Martens and Grosse, 2015) with trust region. Zhao et al. (2021) have demonstrated that this method has a surprising superiority on online 3D packing problems over other model-free DRL algorithms like SAC (Haarnoja et al., 2018).
@@
Specifically, ACKTR runs multiple parallel processes (64 here) to interact with their respective environments and gather samples. The different processes may have different packing time steps $t$ and deal with different packing sequences; the space configuration node number N also changes. To combine these data with irregular shapes into one batch, we fulfill $\textbf{B}_{t}$ and $\textbf{L}_{t}$ to fixed lengths, $80$ and $25\cdot|\textbf{O}|$ respectively, with dummy nodes. The descriptors for dummy nodes are all-zero vectors and have the same size as the internal nodes or the leaf nodes. The relation weight logits $u_{ij}$ of dummy node $j$ to arbitrary node $i$ is replaced with $-inf$ to eliminate these dummy nodes during the feature calculation of GAT. The global context feature $\bar{h}$ is aggregated only on the eligible nodes h: $\bar{h}=\frac{1}{N}\sum_{i=1}^{N}h_{i}$. All space configuration nodes are embedded by GAT as a fully connected graph as Figure 2 (b), without any inner mask operation.
[…]
@@
[…]
@@ Appendix B Leaf Node Expansion Schemes
[…]
We introduce the leaf node expansion schemes adopted in our PCT implementation here. These schemes are used to incrementally calculate new candidate placements introduced by the just-placed item $n_{t}$. A good expansion scheme should reduce the number of solutions to be explored while not missing too many feasible packings. Meanwhile, polynomials computability is also expected. As shown in Figure 17, the policies guided by suitable leaf node expansion schemes outperform the policy trained on a full coordinate (FC) space in the whole training process. We extend three existing heuristic placement rules, which have proven to be both accurate and efficient, to our PCT expansion, i.e. Corner Point, Extreme Point, and Empty Maximal Space. Since all these schemes are related to boundary points of packed items, we combine the start/end points of $n_{t}$ with these boundary points as a superset, namely Event Point.
@@
Martello et al. (2000) first introduce the concept of Corner Point (CP) for their branch-and-bound methods. Given 2D packed items in the $xoy$ plane, the corner points can be found where the envelope of the items in the bin changes from vertical to horizontal, as shown in Figure 18 (a). The past corner points that no longer meet this condition will be deleted.
Extend this 2D situation to 3D cases, the new candidate 3D positions introduced by the just placed item $n_{t}$ are a subset of $\{(p_{n}^{x}+s_{n}^{x},p_{n}^{y},p_{n}^{z}),(p_{n}^{x},p_{n}^{y}+s_{n}^{y},p_{n}^{z}),(p_{n}^{x},p_{n}^{y},p_{n}^{z}+s_{n}^{z})\}$ if the envelope of the corresponding 2D plane, i.e. $xoy,yoz,\text{and }xoz$, is changed by $n_{t}$. The time complexity of finding 3D corner points incrementally is $O(c)$ with an easy-to-maintained bin height map data structure to detect the change of envelope on each plane, $c$ is a constant here.
Figure 18: Full candidate positions generated by different PCT expansion schemes (all in $xoy$ plane). The gray dashed lines are the boundaries of the bin. Circles in (a) and (b) represent corner points and extreme points, respectively. (c) The candidate positions (circles) introduced by different EMSs are rendered with different colors. All intersections of two dashed lines in (d) constitute event points.
@@
Crainic et al. (2008) extend the concept of Corner Point to Extreme Point (EP) and claim their method reaches the best offline performance of that era. Its insight is to provide the means to exploit the free space defined inside a packing by the shapes of the items that already exist. When the current item $n_{t}$ is added, new EPs are incrementally generated by projecting the coordinates $\{(p_{n}^{x}+s_{n}^{x},p_{n}^{y},p_{n}^{z}),(p_{n}^{x},p_{n}^{y}+s_{n}^{y},p_{n}^{z}),(p_{n}^{x},p_{n}^{y},p_{n}^{z}+s_{n}^{y})\}$ on the orthogonal axes, e.g., project $(p_{n}^{x}+s_{n}^{x},p_{n}^{y},p_{n}^{z})$ in the directions of the $y$ and $z$ axes to find intersections with all items lying between item $n_{t}$ and the boundary of the bin. The nearest intersection in the respective direction is an extreme point. Since we stipulate a vertical top-down loading direction, the 3D extreme points in the strict sense may exist a large item blocking the loading direction. So we find the 2D extreme points (see Figure 18 (b)) in the $xoy$ plane and repeat this operation on each distinct $p^{z}$ value (i.e. start/end $z$ coordinate of a packed item) which satisfies $p^{z}_{n}\leq p^{z}\leq p^{z}_{n}+s^{z}_{n}$. The time complexity of this method is $O(m\cdot|\textbf{B}_{2D}|)$, where $\textbf{B}_{2D}$ is the packed items that exist in the corresponding $z$ plane and $m$ is the number of related $z$ scans.
@@
Appendix C More Results
@@
In this section, we report more results of our method. Section C.1 further discusses the generalization ability of our method on disturbed item sampling distributions and unseen items. Section C.2 visualizes packing sequences to analyze model behaviors. Section C.4 reports the running cost of each method. More visualized results are provided in Section D.
C.1 Generalization Performance
@@
The generalization ability of learning-based methods has always been a concern. Here we demonstrate that our method has a good generalization performance on item size distributions different from the training one. We conduct this experiment with continuous solution space. We sample item size $s^{d}$ from normal distributions $N(\mu,\sigma^{2})$ for generating test sequences where $\mu$ and $\sigma$ are the expectation and the standard deviation. Three normal distributions are adopted here, namely $N(0.3,0.1^{2})$, $N(0.1,0.2^{2})$, and $N(0.5,0.2^{2})$, as shown in Figure 20. The larger $\mu$ of the normal distribution, the larger the average size of sampled items. We still control $s^{d}$ within the range of $[0.1,0.5]$. If the sampled item sizes are not within this range, we will resample until they meet the condition.
@@
[…]
@@
We directly transfer our policies trained on $U(0.1,0.5)$ to these new datasets without any fine-tuning. We use the best-performing heuristic method LSAH (Hu et al., 2017) in Section 4.3 as a baseline. The test results are summarized in Table 11. Our method performs well on distributions different from the training one and always surpasses the LSAH method. See Appendix C for more results about the generalization ability of our method on disturbed distributions and unseen items.
We demonstrate that our algorithm has a good generalization performance on disturbed item transitions, i.e., $\mathcal{P}(s_{t+1}|s_{t})$. Generalizing to a new transition is a classic challenge for reinforcement learning (Taylor and Stone, 2009; Liu et al., 2022). We conduct this experiment in the discrete setting where the item set is finite ($|\mathcal{I}|=125$). For each item $n\in\mathcal{I}$, we add a random non-zero disturbance $\delta_{i}$ on its original sample probability $p_{i}$, e.g., $p_{i}=p_{i}\cdot(1-\delta_{i})$. We normalize the disturbed $p_{i}$ as the final item sampling probability. Note that $\delta_{i}$ is fixed during sampling one complete sequence. We test different ranges of $\delta_{i}$ and the results are summarized in Table 12.
Table 12: Transfer the best-performing PCT policies directly to the disturbed item sampling distributions. Dif. means how much the generalization performance drops from the undisturbed case ($\delta_{i}=0$).
[…]
@@
Beyond experiments on generalization to disturbed distributions, we also test our method with unseen items. We conduct this experiment in the discrete setting. We randomly delete 25 items from $\mathcal{I}$ and train PCT policies with $|\mathcal{I}_{sub}|=100$. Then we test the trained policies on full $\mathcal{I}$. See Table 13 for results. Our method still performs well on datasets where unseen items exist regarding all settings.
Table 13: Generalization performance on unseen items. All policies are trained with the EV scheme.
@@
C.2 Understanding of Model Behaviors

### v4

@@ Appendix A Implementation Details
$J(\pi)=E_{\psi\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$ (12)
Where $\gamma\in[0,1]$ is the discount factor, and $\psi=(s_{0},a_{0},s_{1},\ldots)$ is a trajectory sampled based on the policy $\pi$. We extract the feature of state $s_{t}=(\mathcal{T}_{t},n_{t})$ using graph attention networks (Velickovic et al., 2018) for encoding the spatial relations of all space configuration nodes. The context feature is fed to two key components of our pipeline: an actor network and a critic network. The actor network, designed based on a pointer mechanism, weighs the leaf nodes of PCT, which is written as $\pi(a_{t}|s_{t})$. The action $a_{t}$ is an index of selected leaf node $l\in\textbf{L}_{t}$, denoted as $a_{t}=index(l)$. The critic network maps the context feature into a state value prediction $V(s_{t})$, which helps the training of the actor network. The whole network is trained via a composite loss $L=\alpha\cdot L_{actor}+\beta\cdot L_{critic}$ ($\alpha=\beta=1$ in our implementation), which consists of actor loss $L_{actor}$ and critic loss $L_{critic}$. These two loss functions are defined as:
[…]
Where $r_{t}=c_{r}\cdot w_{t}$ is our reward signal, and we set $\gamma$ as 1 since the packing episode is finite. We adopt a step-wise reward $r_{t}=c_{r}\cdot w_{t}$ once $n_{t}$ is inserted into PCT as an internal node successfully. Otherwise, $r_{t}=0$ and the packing episode ends. The choice of item weight $w_{t}$ depends on the packing preferences. In the general sense, we set $w_{t}$ as the volume occupancy $v_{t}=s^{x}_{n}\cdot s^{y}_{n}\cdot s^{z}_{n}$ of $n_{t}$, and the constant $c_{r}$ is $10/(S^{x}\cdot S^{y}\cdot S^{z})$. For online 3D-BPP with additional packing constraints, this weight can be set as $w_{t}=max(0,v_{t}-c\cdot O(s_{t},a_{t}))$. While the term $v_{t}$ ensures that space utilization is still the primary concern, the objective function $O(s_{t},a_{t})$ guides the agent to satisfy additional constraints like isle friendliness and load balancing. We adopt the ACKTR (Wu et al., 2017) method for training our DRL agent. Zhao et al. (2021) have demonstrated that this method has a superiority on online 3D-BPP over other model-free DRL algorithms like SAC (Haarnoja et al., 2018).
@@
Specifically, ACKTR runs multiple parallel processes (64 here) to interact with their respective environments and gather samples. The different processes may have different packing time steps $t$ and deal with different packing sequences; the space configuration node number N also changes. To combine these data with irregular shapes into one batch, we fulfill $\textbf{B}_{t}$ and $\textbf{L}_{t}$ to fixed lengths, $80$ and $25\cdot|\textbf{O}|$ respectively, with dummy nodes. The descriptors for dummy nodes are all-zero vectors and have the same size as the internal nodes or the leaf nodes. The relation weight logits $u_{ij}$ of dummy node $j$ to arbitrary node $i$ is replaced with inf-inf to eliminate these dummy nodes during the feature calculation of GAT. The global context feature $\bar{h}$ is aggregated only on the eligible nodes h: $\bar{h}=\frac{1}{N}\sum_{i=1}^{N}h_{i}$. All space configuration nodes are embedded by GAT as a fully connected graph as Figure 2 (b), without any inner mask operation.
We only provide the packable leaf nodes that satisfy placement constraints for DRL agents. For setting 2setting\,2, we check in advance if a candidate placement satisfies Equation (2) and (3). For setting 1setting\,1 and setting 3setting\,3, where the mass of item $n_{t}$ is $v_{t}$ and $\rho\cdot v_{t}$ respectively, we will additionally check if one placement meets the constraints of packing stability. Benefits from the fast stability estimation method proposed by Zhao et al. (2022b), this pre-checking process can be completed in a very short time, and our DRL agent samples data at a frequency of more than 400 FPS.
@@
Since PCT allows discarding some valid leaf nodes and this will not harm our performance, we randomly intercept a subset $\textbf{L}_{sub_{t}}$ from $\textbf{L}_{t}$ if $|\textbf{L}_{t}|$ exceeds a certain length. Determining the suitable PCT length for different bin configurations is important, and we give our recommendations for finding this hyperparameter. For training, we find that the performance of learned policies is more sensitive to the number of allowed orientations $|\textbf{O}|$. Thus, we set the PCT length as $c\cdot|\textbf{O}|$ where $c$ can be determined by a grid search near $c=25$ for different bin configurations. For our experiments, $c=25$ works quite well. During the test, the PCT length can be different from the training one. We suggest searching for this interception length with a validation dataset via a grid search, which ranges from 50 to 300 with a step length of 10.
@@ Appendix B Leaf Node Expansion Schemes
Figure 17: Learning curves on setting 1setting\,1. A good expansion scheme for PCT reduces the complexity and helps DRL methods for more efficient learning and better performance. EVF means the full EV leaf node set without an interception.
We introduce the leaf node expansion schemes adopted in our PCT implementation here. These schemes are used to incrementally calculate new candidate placements introduced by the just-placed item $n_{t}$. A good expansion scheme should reduce the number of solutions to be explored while not missing too many feasible packings. Meanwhile, polynomials computability is also expected. As shown in Figure 17, the policies guided by suitable leaf node expansion schemes outperform the policy trained on a full coordinate (FC) space in the whole training process. We extend three existing heuristic placement rules, which have proven to be both accurate and efficient, to our PCT expansion, i.e., Corner Point, Extreme Point, and Empty Maximal Space. Since all these schemes are related to boundary points of packed items, we combine the start/end points of $n_{t}$ with these boundary points as a superset, namely Event Point.
@@
Martello et al. (2000) first introduce the concept of Corner Point (CP) for their branch-and-bound methods. Given 2D packed items in the xoyxoy plane, the corner points can be found where the envelope of the items in the bin changes from vertical to horizontal, as shown in Figure 18 (a). The past corner points that no longer meet this condition will be deleted.
Extend this 2D situation to 3D cases, the new candidate 3D positions introduced by the just placed item $n_{t}$ are a subset of $\{(p_{n}^{x}+s_{n}^{x},p_{n}^{y},p_{n}^{z}),(p_{n}^{x},p_{n}^{y}+s_{n}^{y},p_{n}^{z}),(p_{n}^{x},p_{n}^{y},p_{n}^{z}+s_{n}^{z})\}$ if the envelope of the corresponding 2D plane, i.e. $xoy,yoz,\text{and }xoz$, is changed by $n_{t}$. The time complexity of finding 3D corner points incrementally is $O(c)$ with an easy-to-maintained bin height map data structure to detect the change of envelope on each plane, where $c$ is a constant.
Figure 18: Full candidate positions generated by different PCT expansion schemes (all in xoyxoy plane). The gray dashed lines are the boundaries of the bin. Circles in (a) and (b) represent corner points and extreme points, respectively. (c) The candidate positions (circles) introduced by different EMSs are rendered with different colors. All intersections of two dashed lines in (d) constitute event points.
@@
Crainic et al. (2008) extend the concept of Corner Point to Extreme Point (EP) and claim their method reaches the best offline performance of that era. Its insight is to provide the means to exploit the free space defined inside a packing by the shapes of the items that already exist. When the current item $n_{t}$ is added, new EPs are incrementally generated by projecting the coordinates $\{(p_{n}^{x}+s_{n}^{x},p_{n}^{y},p_{n}^{z}),(p_{n}^{x},p_{n}^{y}+s_{n}^{y},p_{n}^{z}),(p_{n}^{x},p_{n}^{y},p_{n}^{z}+s_{n}^{y})\}$ on the orthogonal axes, e.g., project $(p_{n}^{x}+s_{n}^{x},p_{n}^{y},p_{n}^{z})$ in the directions of the $y$ and $z$ axes to find intersections with all items lying between item $n_{t}$ and the boundary of the bin. The nearest intersection in the respective direction is an extreme point. Since we stipulate a top-down loading direction, the 3D extreme points in the strict sense may exist a large item blocking the loading direction. So we find the 2D extreme points (see Figure 18 (b)) in the xoyxoy plane and repeat this operation on each distinct $p^{z}$ value (i.e. start/end $z$ coordinate of a packed item) which satisfies $p^{z}_{n}\leq p^{z}\leq p^{z}_{n}+s^{z}_{n}$. The time complexity of this method is $O(m\cdot|\textbf{B}_{2D}|)$, where $\textbf{B}_{2D}$ is the packed items that exist in the corresponding $z$ plane and $m$ is the number of related $z$ scans.
@@
Appendix C Theoretical Analysis of PCT
Although PCT is guided by heuristic rules that also restrict the solution space, we provide a theoretical proof that certain heuristic-generated candidates are locally optimal. Reinforcement learning then builds upon these local optima to ultimately make a global decision.
Given an item $n$, the goal is to identify a finite set of candidate placements where $n$ is tightly positioned against obstacles (i.e., the container or packed items), such that it permits no further movement and leaves maximal space for future placements. Since we adopt a top-down placement manner that determines item positions on the xoyxoy plane, the 3D packing can be reduced to a 2.5D formulation, i.e., analyzing optimal placement on each horizontal 2D plane with the same $z$ value. For an item $n$ with a known orientation, we denote its horizontal projection $G\subset\mathbb{R}^{2}$, which is a 2D polygon. Accordingly, we extract 2D No-Fit Polygons (NFP) $E$ from each $z$ plane. The NFP (Burke et al., 2007) is a geometric representation of all relative positions where one polygon can touch another without overlapping, as shown in Figure 20 (a). Let $O\subset\mathbb{R}^{2}$ denote the 2D projection of packed items; we can get NFP $E$ with $E\oplus G\subset C$ and $E\oplus G\cap O=\emptyset$, where $\oplus$ is the Minkowski sum (Mark et al., 2008). A 2D theorem can be established to show that the convex vertices of $E$ are local optimum for tight packing:
Theorem 1.
For a boundary point $p\in\partial E$, if $p$ is a convex vertex, then $p=\text{argmax}_{q\in\mathcal{N}(p)}\psi(q)$ for an open neighborhood $\mathcal{N}(p)$ of $p$ and a tightness measure $\psi(\cdot)$.
Figure 20: Select a reference point $p$ from the NFP $E$ (a), its corresponding placement is $\{p\}\oplus G$ (b). (c): In direction $d$, the projection $e_{q}=e(G,q,d)$ of $\{q\}\oplus G$ exceeds $e_{p}$ of $\{p\}\oplus G$, indicating $\{q\}\oplus n$ is a more extreme placement in $d$.
We formally define our tightness measure $\psi(\cdot)$ and provide a proof of Theorem 1. The placement of item $n$ refers to selecting a reference point $p$ from $E$ with $\{p\}\oplus G$ being the corresponding placement, as illustrated in Figure 20 (b). We prefer reference points that result in $n$ being tightly packed against obstacles. However, quantifying this tightness is nontrivial. Mark et al. (2008) introduce the notion of extremeness for polygons: a polygon is considered more extreme in direction $d$ if its furthest point along $d$ lies farther than that of another. Inspired by this, we define the tightness measure as $e(G,p,d)=\max_{g\in\{p\}\oplus G}d^{T}g$, which is the extreme value of $G$’s projection for the packing task. Placement at $p$ is more extreme than at $q$ if $e(G,p,d)>e(G,q,d)$. When $G$ is tightly packed against obstacles along $d$, $e(G,p,d)$ achieves a local maximum, as illustrated in Figure 20 (c). This leads to an intuitive tightness criterion:
$p=\text{argmax}_{q\in\mathcal{N}(p)}\;e(G,q,d)$ (14)
for some open neighborhood $\mathcal{N}(p)\subset E$ and direction $d\in\mathbb{R}^{2}$. The following property arises directly from the Minkowski sum, establishing a geometric connection between tight packing and feasible reference points in $E$:
Lemma 1.
If Equation (14) holds for some open neighborhood $\mathcal{N}(p)$, then:
$p=\text{argmax}_{q\in\mathcal{N}(p)}d^{T}q$ (15)
In other words, if $d^{T}p$ is a local maximum, i.e., $d^{T}p\geq d^{T}q$ for all $q\in\mathcal{N}(p)$, then $p$ represents the most extreme placement of $G$ in direction $d$ within its neighborhood. With this formulation, we can assess the packing tightness of point $p$ by evaluating the range of directions $d$ for which $p$ satisfies Equation (15). This range corresponds to the spanning angle of the 2D normal cone (Boyd et al., 2004). The normal cone of a set $C$ at a boundary point $x_{0}$ is the set of all vectors $y$ such that $y^{T}(x-x_{0})\leq 0$ for all $x\in C$. Based on this, we define the tightness measure:
Lemma 2.
For a convex polygonal set $E$ and $p\in\partial E$, the tightness measure $\psi(p)$ is defined as the spanning angle of the normal cone at $p$, given by $\psi(p)\triangleq\pi-\theta$, where $\theta$ is the interior angle of $E$ at $p$.
Figure 21: (a): For a boundary point $p$ on a convex cone $E$, any direction $d$ in its normal cone satisfies $d^{T}p\geq d^{T}q$, where $q\in E$. (b): If $p$ lies on a concave segment, then for $\forall d\in\mathbb{R}^{2}$, there exists $q\in E$ that the angle $\vec{pq},d><\pi/2$, i.e., $d^{T}p<d^{T}q$.
Figure 21 (a) illustrates an example of a normal cone. For concave polygons, points on concave boundaries have no normal cone, as shown in Figure 21 (b). We assign $\psi(p)=0$ in these cases. Using $\psi(p)$ as a tightness metric, we compare different candidate placements. The proof for Theorem 1 is obvious:
Proof.
Normal cones exist only at boundary points, not interior ones. Consider a neighborhood $\mathcal{N}(p)$ that includes $p$ and its two neighboring half-open edges. Within $\mathcal{N}(p)$, point $p$ is convex with $\psi(p)=\pi-\theta$. All other points $q$ are with $\psi(q)=0$. ∎
Figure 22: (a): Candidates that are locally optimal under the tightness measure $\psi$ are covered by the EMSs, with (b) and (c) illustrating their FLB coordinates and final placements, respectively. $\mathcal{N}(p_{1})$ and $\mathcal{N}(p_{2})$ are the neighborhoods for convex vertices $p_{1}$ and $p_{2}$.
This proof provides a geometric interpretation of why PCT performs well under heuristic-based leaf node expansion schemes. Some schemes, such as EMS and EV, operate along the boundaries of already packed items in the $d\in\{x,y\}$ directions. The intersecting boundaries cover the convex vertices of the NFP, thereby preserving local optimality while pruning much of the solution space to reduce unnecessary exploration. As illustrated in Figure 22 (a), each open circle represents an EMS candidate that can be transformed into a locally optimal FLB candidate. Examples of such transformed FLB candidate (filled circle) and their corresponding placements are exhibited in Figure 22 (b) and (c), respectively. The DRL policy then selects global placements from these high-quality local candidates, enhancing both learning efficiency and performance.
Appendix D More Results
@@
In this section, we report more results of our method. Section D.1 further discusses the generalization ability of our method on disturbed item sampling distributions and unseen items. Section D.2 visualizes packing sequences to analyze model behaviors. Section D.4 reports the running cost of each method.
D.1 Generalization Performance
@@
The generalization ability of learning-based methods has always been a concern. Here we demonstrate that our method has a good generalization performance on item size distributions different from the training one. We conduct this experiment with continuous solution space. We sample item size $s^{d}$ from normal distributions $\mathcal{N}(\mu,\sigma^{2})$ for generating test sequences where $\mu$ and $\sigma$ are the expectation and the standard deviation. Three normal distributions are adopted here, namely $\mathcal{N}(0.3,0.1^{2})$, $\mathcal{N}(0.1,0.2^{2})$, and $\mathcal{N}(0.5,0.2^{2})$, as shown in Figure 23. The larger $\mu$ of the normal distribution, the larger the average size of sampled items. We still control $s^{d}$ within the range of $[0.1,0.5]$. If the sampled item sizes are not within this range, we will resample until they meet the condition.
@@
Test Distribution Method Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
@@
We directly transfer our policies trained on $\mathcal{U}(0.1,0.5)$ to these new datasets without any fine-tuning. We use the best-performing heuristic method LSAH (Hu et al., 2017) in Section 4.3 as a baseline. The test results are summarized in Table 12. Our method performs well on distributions different from the training one and always surpasses the LSAH method.
We demonstrate that our algorithm has a good generalization performance on disturbed item transitions, i.e., $\mathcal{P}(s_{t+1}|s_{t})$. Generalizing to a new transition is a classic challenge for reinforcement learning (Taylor and Stone, 2009). We conduct this experiment in the discrete setting where the item set is finite ($|\mathcal{I}|=125$). For each item $n\in\mathcal{I}$, we add a random non-zero disturbance $\delta_{i}$ on its original sample probability $p_{i}$, e.g., $p_{i}=p_{i}\cdot(1-\delta_{i})$. We normalize the disturbed $p_{i}$ as the final item sampling probability. Note that $\delta_{i}$ is fixed during sampling one complete sequence. We test different ranges of $\delta_{i}$ and the results are summarized in Table 13.
Table 13: Transfer the best-performing PCT policies directly to the disturbed item sampling distributions. “Dif.” means how much the generalization performance drops from the undisturbed case ($\delta_{i}=0$).
Disturbance Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
@@
Beyond experiments on generalization to disturbed distributions, we also test our method with unseen items. We conduct this experiment in the discrete setting. We randomly delete 25 items from $\mathcal{I}$ and train PCT policies with $|\mathcal{I}_{sub}|=100$. Then we test the trained policies on full $\mathcal{I}$. See Table 14 for results. Our method still performs well on datasets where unseen items exist in all settings.
Table 14: Generalization performance on unseen items.
Train Test Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
@@
D.2 Understanding of Model Behaviors